receivables transactions.

E8.4  (LO 1, 2), AP Assume the following information for Larry Corp.

Accounts receivable (beginning balance)

$142,000

Allowance for doubtful accounts (beginning balance)

11,360

Net credit sales

945,000

Collections

910,000

Write-offs of accounts receivable

5,200

Collections of accounts previously written off

1,900

Uncollectible accounts are expected to be 8% of the ending balance in accounts receivable.

Instructions

a.  

Prepare the entries to record sales and collections during the period.

b.  

Prepare the entry to record the write-off of uncollectible accounts during the period.

c.  

Prepare the entries to record the recovery of the uncollectible account during the period.

d.  

Determine the ending balance in Accounts Receivable and the unadjusted balance in Allowance for Doubtful Accounts.

e.  

Prepare the entry to record bad debt expense for the period.

f.  

Determine the ending (adjusted) balance in Allowance for Doubtful Accounts.
Answer #1

Journal entry

No Account and explanation Debit Credit
a Account receivable 945000
Sales revenue 945000
(To record sales)
Cash 910000
Account receivable 910000
(To record collection)
b Allowance for doubtful accounts 5200
Account receivable 5200
(To record write off)
c Account receivable 1900
Allowance for doubtful account 1900
(To record reinstate)
Cash 1900
Account receivable 1900
(To record collection)

d) Ending balance of account receivable = 142000+945000-910000-5200 = 171800

Unadjusted balance for allowance for doubtful accounts = 11360+1900-5200 = 8060

e) Bad debt expense = (171800*8%-8060) = 5684

f) Ending balance in allowance for doubtful accounts = 13744